OpenStudy (anonymous):

in a candy bar,165 calories are from sugar.there are 250 calories in the bar.what percent of the calories are from sugar?

OpenStudy (anonymous):

You divide 165 by 250 to .66 and then you move the decimal two places to the right to get 66% OR you can set up an cross multiplying equation to do it. \[\frac{ 165 }{ 250 } \times \frac { x }{ 100} \] So multiply 165 by 100 and then divide by 250 to get 66%.